### v1

Abstract.
For an action of a discrete group $\Gamma$ on a set $X$, we show that the Schreier graph on $X$ is property A if and only if the permutation representation on $\ell_{2}X$ generates an exact $\operatorname{\mathrm{C}^{*}}$-algebra. This is well known in the case of the left regular action on $X=\Gamma$. This also generalizes Sako’s theorem, which states that exactness of the uniform Roe algebra $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X)$ characterizes property A of $X$ when $X$ is uniformly locally finite.
Key words and phrases:
property A, exactness, Schreier graph
1991 Mathematics Subject Classification:
46L05, 51F99
1. Introduction
Guoliang Yu introduced Property A as a kind of amenable-type condition on metric spaces in [Yu00]. It is characterized in terms of operator algebras by Skandalis, Tu, and Yu in [STY02, Theorem 5.3], that is property A of a metric space X with uniformly locally finite (or ulf in short) is equivalent to nuclearity of the uniform Roe algebra $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X)$. Moreover, by Sako’s remarkable result ([Sak20, Theorem 1.1]), exactness and local reflexivity of $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X)$ also characterize property A under ulf. These results hold as well when $X$ is a ulf coarse space. See [Sak13a].
For a finitely generated group $\Gamma$, we equip $\Gamma$ with the word metric (or generally the coarse structure) and regard it as a ulf metric space $\left\lvert\Gamma\right\rvert$. In the case of discrete groups, it is well-known that property A is equivalent to $\operatorname{\mathrm{C}^{*}}$-exactness, i.e. reduced group $\operatorname{\mathrm{C}^{*}}$-algebra $\mathrm{C}^{*}_{\mathrm{red}}(\Gamma)$ being exact. Kirchberg and Wassermann introduce the notion of exact groups in terms of reduced crossed products and characterize by exactness of $\mathrm{C}^{*}_{\mathrm{red}}(\Gamma)$ ([KW99, Theorem 5.2]). Ozawa proves that this is equivalent to nuclearity of $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}\left\lvert\Gamma\right\rvert$ in [Oza00, Theorem 3], which is identical with property A of $\left\lvert\Gamma\right\rvert$ ([HR00, Theorem 1.1][ADR00]).
The purpose of this paper is to combine and generalize these results.
Theorem A (Main Theorem).
For a discrete group $\Gamma$ and an action on a set $X$, the Schreier graph $\left\lvert\Gamma\curvearrowright X\right\rvert$ is property A if and only if the permutation representation $\lambda_{X}:\Gamma\to\mathbb{B}(\ell_{2}X)$ generates the exact $\operatorname{\mathrm{C}^{*}}$-algebra $\operatorname{\mathrm{C}^{*}}(\lambda_{X}(\Gamma))$.
Moreover, for a set $\Gamma$ of partial translations on a set $X$, $\Gamma$ generates the ulf coarse structure on $X$ of property A if and only if $\Gamma$ generates the exact $\operatorname{\mathrm{C}^{*}}$-subalgebra $\operatorname{\mathrm{C}^{*}}(\Gamma)$ of $\mathbb{B}(\ell_{2}X)$.
We prove this partial translation version in Section 4 by the $2\times 2$ trick. We remark that our theorem contains Sako’s result about exactness of $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X)$ because every ulf coarse space is generated by partial translations. Although it is not apparent whether local reflexivity of $\operatorname{\mathrm{C}^{*}}(\lambda_{X}(G))$ characterizes property A or not, we can give another proof of Sako’s result about local reflexitivity of $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X)$ by using the lemma 9 in the “generalized box space” setting of [Sak20]. We prove the main theorem in the same spirit as Sako’s paper [Sak20], which discovers the importance of a weighted trace and treats it like an amenable trace. It does not seem that CPAP rephrasing works well, in contrast to the group case.
Corollary B.
Let $X$ be an undirected graph of uniformly bounded degree in which loops and multiple edges are allowed. For any partial translations $v_{1},\dots,v_{l}$ which generate $X$ as a graph, exactness of $\operatorname{\mathrm{C}^{*}}(v_{1},\dots,v_{l})$ only depends on the coarse structure of $X$.
We remark that giving $v_{1},\dots,v_{l}$ is equivalent to giving an orientation and an $l$-edge coloring on $X$ so that two edges exiting or entering the same vertex have different colors.
Acknowledgements
The author would like to thank Narutaka Ozawa for his support and helpful comments as my supervisor.
2. Preliminaries
Although we treat general coarse spaces, we reduce everything to the graph case by regarding general coarse structures as the directed increasing union of graphs. We give a brief introduction to ulf coarse spaces in Section 4. For more details, see standard references on coarse geometry ([Roe03][NY12]).
2.1. Notations
For an action of a finitely generated group $\Gamma$ on a set $X$, define the Schreier graph $\left\lvert\Gamma\curvearrowright X\right\rvert$ by the graph $(X,E)$, where $E:=\{\,(sx,x)\in X\times X\mid x\in X,s\in S\,\}$ and $S\subset\Gamma$ is a symmetric finite generator set. This coincides with the Cayley graph when the action is the left regular action $\Gamma\curvearrowright\Gamma$.
Let $(X,E)$ be a (possiblly non-connected) undirected graph and $d$ be its graph metric ($d$ takes a value in $[0,\infty]$). We assume that $\sup_{x\in X}\left\lvert E\cap(X\times\{x\})\right\rvert<\infty$, which is called uniformly locally finite, or ulf in short (a.k.a. of uniformly bounded degree, of bounded geometry). For $R>0$, we denote by
$\mathbb{C}_{\mathrm{u}}^{\leq R}(X):=\left\{\,a\in\mathbb{B}(\ell_{2}X)\mathrel{}\middle|\mathrel{}a_{x,y}=0\ \text{if}\ d(x,y)>R\,\right\}$
and define the uniform Roe algebra $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X)$ by norm closure of $\bigcup_{R>0}\mathbb{C}_{\mathrm{u}}^{\leq R}(X)$.
For $S>0,x\in X$, we recall the compression $\Phi_{S}$ in [Sak14]
$\Phi_{S}:\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X)\to\prod_{x\in X}\mathbb{B}(\ell_{2}\operatorname{Ball}(x,S)),\ \Phi_{S}(a):=\left(P_{\operatorname{Ball}(x,S)}aP^{*}_{\operatorname{Ball}(x,S)}\right)_{x},$
where $P_{V}$ stands for the orthogonal projection from $\ell_{2}X$ onto $\ell_{2}V$ for $V\subset X$ and $\operatorname{Ball}(x,S):=\{\,y\in X\mid d(x,y)\leq S\,\}$.
For a Hilbert space $\mathcal{H}$, denote $\mathbb{B}(\mathcal{H}),\mathbb{K}(\mathcal{H}),\operatorname{\mathrm{HS}}(\mathcal{H})$ by the $\operatorname{\mathrm{C}^{*}}$-algebra of bounded operators, the $\operatorname{\mathrm{C}^{*}}$-algebra of compact operators, and the Hilbert space of Hilbert-Schmidt operators, respectively. When a $\operatorname{\mathrm{C}^{*}}$-algebra $A\subset\mathbb{B}(\mathcal{H})$ is concretely represented, we define the vector state $\varphi_{\xi}(a):=\mathopen{\langle}\xi,a\xi\mathclose{\rangle}$ for a unit vector $\xi\in\mathcal{H}$. Let $\operatorname{Prob}(X)\subset\ell_{1}X$ be the space of probability measures on $X$.
2.2. Property A
Property A has many characterizations, like amenability in the group theory. We refer to some of them except the original definition in [Yu00].
Definition 1 (See [CTWY08, Section 2]).
A metric space $X$ with bounded geometry is said to have the operator norm localization property or ONL in short, if for every $\varepsilon,R>0$ there exists $S>0$ such that $\left\lVert\Phi_{S}(a)\right\rVert\geq(1-\varepsilon)\left\lVert a\right\rVert$ for every $a\in\mathbb{C}^{\leq R}_{\mathrm{u}}(X)$.
Theorem 2 (See [STY02, Theorem 5.3][Sak14, Theorem 4.1]).
For a metric space $X$ of bounded geometry, the following are all equivalent.
  • $X$ has property A.
  • $X$ has the ONL.
  • $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X)$ is nuclear.
This theorem also holds when $X$ is a ulf coarse space. See [Sak13a].
Lemma 3.
Let $(X,E)$ be a ulf graph and $\varepsilon>0$. Then there exists $L>1$ such that for every $\eta\in\operatorname{Prob}(X)$ with a finite support, there exists $\eta^{\prime}\in\operatorname{Prob}(X)$ satisfying that
  • $\left\lVert\eta-\eta^{\prime}\right\rVert_{1}\leq\varepsilon$,
  • $L^{-1}\eta^{\prime}(x)\leq\eta^{\prime}(y)\leq L\eta^{\prime}(x)$ for every $(x,y)\in E$,
  • $\left\lVert P_{V}\eta^{\prime}\right\rVert\leq\varepsilon^{\operatorname{dist}(V,\operatorname{\mathop{supp}}\eta)}$ for $V\subset X$, where $P_{V}$ stands the canonical projection from $\ell_{1}X$ onto $\ell_{1}V$.
Although this lemma was essentially proven in [Sak20, Theorem 3.17 (4),(5)], we give its proof for the reader’s convenience.
Proof.
Define $M:=\sup_{x\in X}\left\lvert E\cap(\{x\}\times X)\right\rvert<\infty$ and $L:=\max\{1,M(1+\varepsilon^{-1})\}$.
$\eta^{\prime}(x):=\sum_{w}L^{-d(x,w)}\eta(w)$
satisfies the three conditions because
  • $\eta^{\prime}\geq\eta$ and $\sum_{x}L^{-d(x,w)}\leq 1+\epsilon$ for every $w\in X$,
  • $d(x,w)+1\geq d(y,w)\geq d(x,w)-1$ for every $(x,y)\in E$,
  • $\sum_{d(x,w)\geq R}L^{-d(x,w)}\leq\epsilon^{R}(1+\epsilon)^{-R+1}$ for every $w\in X,R>0$.
Then we adjust $\eta^{\prime}$ to have norm 1 and $L$ to be a bit larger. ∎
Let $(X,E)$ be a ulf graph without property A. Then, by [RW14, Lemma 4.2] or [Sak20, Lemma 2.8], there exists $R>0$ and a sequence $b_{n}\in\mathbb{C}^{\leq R}_{\mathrm{u}}(X)$ of positive elements with norm 1 satisfying that
  • $\limsup_{n}\left\lVert\Phi_{S}(b_{n})\right\rVert\leq 0.1$ for every $S>0$,
  • there exist sparse subsets $V_{n}\Subset X$ such that $b_{n}=P_{V_{n}}b_{n}P_{V_{n}}$,
where we call $\{V_{n}\}_{n}$ sparse if $\operatorname{dist}(V_{n},V_{m})\to\infty$ as $n,m\to\infty$. Take $\xi_{n}\in\ell_{2}V_{n}$ as $b_{n}\xi_{n}=\xi_{n}$. By conjugation of $b_{n}$ by a unitary in $\ell_{\infty}X\subset\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X)$, we may assume $\xi_{n}$ to be entrywise non-negative. Using the previous lemma to $(\xi_{n})^{2}$, we take $\xi^{\prime}_{n}\in\ell_{2}X$ which converges to 0 in weak topology because of the exponential decay of $(\xi^{\prime}_{n})^{2}$ and sparse supports $\operatorname{\mathop{supp}}\xi_{n}$. Therefore, we obtain the following:
Lemma 4.
Let $(X,E)$ be a ulf graph without property A and $\varepsilon>0$. Then there exists $L>0$, a sequence $b_{n}\in\mathbb{C}^{\leq R}_{\mathrm{u}}(X)$ of positive elements with norm 1 and a sequence $\xi_{n}\in\ell_{2}X$ of non-negative (in the sense of entrywise) unit vectors which converges to 0 in weak topology satisfying that
  • $\limsup_{n}\left\lVert\Phi_{S}(b_{n})\right\rVert\leq 0.1$ for every $S>0$,
  • $\left\lVert b_{n}\xi_{n}-\xi_{n}\right\rVert\leq\varepsilon$,
  • $L^{-1}\xi_{n}(x)^{2}\leq\xi_{n}(y)^{2}\leq L\xi_{n}(x)^{2}$ for every $(x,y)\in E$.
2.3. Examples
Example 5 (Box space).
[…]
First, we consider $\operatorname{\mathrm{C}^{*}_{\mathrm{RF}}}(\Gamma):=\overline{\mathbb{C}[\Gamma]}\subset\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(\Box\Gamma)$. By residually finiteness, the canonical trace $\tau$ on $\operatorname{\mathrm{C}^{*}_{\mathrm{RF}}}(\Gamma)$ is an amenable trace. Since $\operatorname{\mathrm{C}^{*}_{\mathrm{RF}}}(\Gamma)$ is exact, $\tau$ on $\operatorname{\mathrm{C}^{*}_{\mathrm{red}}}(\Gamma)$ is also amenable, which is equivalent to the amenability of $\Gamma$ by using the characterization of an amenable trace in [BO08, Theorem 6.2.7]. Along this line, Sako proved that exactness of $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X)$ implies property A of $X$ in [Sak20].
In this case, our method is to use directly the min-continuity of
$\tau:\operatorname{\mathrm{C}^{*}_{\mathrm{red}}}(\Gamma)\otimes{\operatorname{\mathrm{C}^{*}_{\mathrm{RF}}}(\Gamma)}^{\mathrm{op}}\to\mathbb{C},\ \ \tau(a\otimes{b}^{\mathrm{op}}):=\tau(ab)$
by Theorem 8. Let $\zeta^{i}\in\ell_{2}\Gamma\otimes\overline{\ell_{2}X}\otimes\ell_{2}$ be vectors that approximate $\tau$ as the vector states. Taking the norm, we reduce $\zeta^{i}$ to $\eta^{i}\in\ell_{2}\Gamma$, which are approximately invariant vectors of $\Gamma$ by Lemma 9.
Example 6.
In this example, we show that nuclearity of $\operatorname{\mathrm{C}^{*}}(v_{1},\dots,v_{l})$ is dependent on the $l$-edge coloring and orientation. Let $X$ be the Cayley graph of $\mathbb{F}_{2}$ (or the regular tree of degree $4$) and $Y$ be the rooted regular tree of out-degree $3$. Note that $X$ is the graph of two copies of $Y$ connected by an edge $v$ with their roots and $\ell_{2}Y$ is the full Fock space over $\mathbb{C}^{3}$ canonically. $\mathbb{F}_{2}\curvearrowright X$ generates the exact (but not nuclear) $\operatorname{\mathrm{C}^{*}}$-algebra $\operatorname{\mathrm{C}^{*}_{\mathrm{red}}}(\mathbb{F}_{2})$. Since the creation operators $T_{1},T_{2},T_{3}$ on $\ell_{2}Y$ generate the Cuntz-Toeplitz algebra $\mathcal{T}_{3}$, the transposition $V$ by $v$ and $T_{i}\otimes 1$ ($i=1,2,3$) on $\ell_{2}X=\ell_{2}Y\otimes\mathbb{C}^{2}$ generate
$\left\{\begin{pmatrix}a&b\\ b&a\end{pmatrix}\mathrel{}\middle|\mathrel{}a\in\mathcal{T}_{3},\ b\in\mathbb{K}\right\}\ \subset\mathbb{B}(\ell_{2}Y\otimes\mathbb{C}^{2}),$
which is nuclear because this is an extension of $\mathcal{O}_{2}$ by $\mathbb{K}\otimes\mathbb{M}_{2}$.
3. Proof of Main Theorem
To prove the main theorem, we mimic the proof in [Sak20]. Consider $X\times\mathbb{N}$, the disjoint union of infinite copies of $X$, and treat this like a “generalized box space” in [Sak13b]. Then, define the “left regular representation” that appeared in [Sak20] as “$(\pi_{\infty},\mathcal{H}_{\infty},\Xi_{\infty})$”, which almost factors through the compression $\Phi_{S}$. We use Hahn-Banach theorem to treat the min-continuity directly.
Denote $\varphi_{\xi}$ by the vector state of a unit vector $\xi$. For vectors $\xi,\xi^{\prime}$ (or just complex numbers), we write $\xi\approx_{\varepsilon}\xi^{\prime}$ when $\left\lVert\xi-\xi^{\prime}\right\rVert\leq\varepsilon$.
Theorem A (Main Theorem).
For a discrete group $\Gamma$ and an action on a set $X$, its Schreier graph $\left\lvert\Gamma\curvearrowright X\right\rvert$ is property A if and only if the permutation representation $\lambda_{X}:\Gamma\to\mathbb{B}(\ell_{2}X)$ generates exact $\operatorname{\mathrm{C}^{*}}$-algebra $\operatorname{\mathrm{C}^{*}}(\lambda_{X}(\Gamma))$.
First, property A of $\left\lvert\Gamma\curvearrowright X\right\rvert$ implies exactness of $\operatorname{\mathrm{C}^{*}}(\lambda_{X}(\Gamma))$ since it is a subalgebra of the nuclear $\operatorname{\mathrm{C}^{*}}$-algebra $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}\left\lvert\Gamma\curvearrowright X\right\rvert$.
We prove this theorem when $\Gamma$ is finitely generated. Then the general case follows immediately because exactness passes to a subalgebra $\operatorname{\mathrm{C}^{*}}(\lambda_{X}(\Lambda))\subset\operatorname{\mathrm{C}^{*}}(\lambda_{X}(\Gamma))$ and nuclearity is preserved under increasing union $\operatorname{\mathrm{C}^{*}}(\ell_{\infty}X,\lambda_{X}(\Gamma))=\lim_{\Lambda\in\mathcal{F}}\operatorname{\mathrm{C}^{*}}(\ell_{\infty}X,\lambda_{X}(\Lambda))$, where $\mathcal{F}$ is the directed family of finitely generated subgroups of $\Gamma$. Note that nuclearity of $\operatorname{\mathrm{C}^{*}}(\ell_{\infty}X,\lambda_{X}(\Lambda))=\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}\left\lvert\Lambda\curvearrowright X\right\rvert$ is equivalent to property A of the Schreier graph $\left\lvert\Lambda\curvearrowright X\right\rvert$ w.r.t. the restricted action of $\Lambda$ on $X$.
We assume that $\operatorname{\mathrm{C}^{*}}(\lambda_{X}(\Gamma))$ is exact and its Schreier graph $\left\lvert\Gamma\curvearrowright X\right\rvert$ is not property A and $\Gamma$ is finitely generated. Therefore, we treat the Schreier graph $\left\lvert\Gamma\curvearrowright X\right\rvert$ as a graph. Take $\xi_{n}\in\ell_{2}X$ and $b_{n}\in\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X)$ as in the previous lemma. For each $\gamma\in\Gamma$, define $h^{\gamma}_{n}\in\ell_{\infty}X$ by $h^{\gamma}_{n}(x)\xi_{n}(\gamma^{-1}x)=\xi_{n}(x)$ and $h^{\gamma}_{n}(x):=1$ if $\xi_{n}(x)=0$ , then $h^{\gamma}:=\oplus_{n}h^{\gamma}_{n}\in\ell_{\infty}(X\times\mathbb{N})$ is invertible.
Consider the left-right representation $\pi^{\prime}$ of $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X)\otimes{\operatorname{\mathrm{C}^{*}}(\lambda_{X}(\Gamma))}^{\mathrm{op}}$ on $\operatorname{\mathrm{HS}}(\ell_{2}X)$, that is $\pi^{\prime}(a\otimes{b}^{\mathrm{op}})\Omega:=a\Omega b$ for $\Omega\in\operatorname{\mathrm{HS}}(\ell_{2}X)$. This is min-countinuous because $\operatorname{\mathrm{HS}}(\ell_{2}X)=\ell_{2}X\otimes\overline{\ell_{2}X}$. Let $\mathcal{U}$ be a (non-principle) ultrafilter on $\mathbb{N}$. Taking the ultrapower, we obtain the representation $\prod_{\mathcal{U}}\pi^{\prime}$ of $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}\left\lvert\Gamma\curvearrowright(X\times\mathbb{N})\right\rvert\otimes{\operatorname{\mathrm{C}^{*}}(\lambda_{X}(\Gamma))}^{\mathrm{op}}$ on $\prod_{\mathcal{U}}\operatorname{\mathrm{HS}}(\ell_{2}X)$. Denote $\lim_{\mathcal{U}}:\prod_{\mathcal{U}}\operatorname{\mathrm{HS}}(\ell_{2}X)\to\operatorname{\mathrm{HS}}(\ell_{2}X)$ by the map taking the weak limit via $\mathcal{U}$. Let
$\mathcal{H}_{0}:=\left(\bigcup_{R>0}\prod_{\mathcal{U}}\operatorname{\mathrm{HS}}^{\leq R}(\ell_{2}X)\right)\cap\mathrm{Ker}\left(\lim_{\mathcal{U}}\right)\ \subset\ \prod_{\mathcal{U}}\operatorname{\mathrm{HS}}(\ell_{2}X)$
be the subspace of finite propagation Hilbert-Schmidt operators, where $\operatorname{\mathrm{HS}}^{\leq R}(\ell_{2}X):=\operatorname{\mathrm{HS}}(\ell_{2}X)\cap\mathbb{C}^{\leq R}_{\mathrm{u}}(X)$. Denote $\mathcal{H}$ by its norm closure.
Lemma 7.
$\mathcal{H}$ is a invariant subspace of $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}\left\lvert\Gamma\curvearrowright(X\times\mathbb{N})\right\rvert\otimes{\operatorname{\mathrm{C}^{*}}(\lambda_{X}(\Gamma))}^{\mathrm{op}}$.
Proof.
Since the finite propagation-ness is preserved under the multiplication of finite propagation operators $\mathbb{C}^{\leq R}_{\mathrm{u}}\left\lvert\Gamma\curvearrowright(X\times\mathbb{N})\right\rvert$, we only show that $\mathrm{Ker}(\lim_{\mathcal{U}})$ is invariant. It suffices to show that $\mathrm{Ker}(\lim_{\mathcal{U}})$ is invariant under the multiplication of $\ell_{\infty}(X\times\mathbb{N})$ because $\mathbb{B}(\operatorname{\mathrm{HS}}(\ell_{2}X))$ commutes with $\lim_{\mathcal{U}}$. To see this, we check that $\{\,f\Omega\in\operatorname{\mathrm{HS}}(\ell_{2}X)\mid f\in\ell_{\infty}X,\ \left\lVert f\right\rVert\leq 1\,\}$ is compact in the norm topology for $\Omega\in\operatorname{\mathrm{HS}}(\ell_{2}X)$. ∎
Since $\mathcal{H}$ is a invariant subspace of $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}\left\lvert\Gamma\curvearrowright(X\times\mathbb{N})\right\rvert\otimes{\operatorname{\mathrm{C}^{*}}(\lambda_{X}(\Gamma))}^{\mathrm{op}}$, we denote $\pi$ by the restriction of $\prod_{\mathcal{U}}\pi^{\prime}$ on $\mathcal{H}$. $\Xi_{n}:=\operatorname{diag}(\xi_{n})\in\operatorname{\mathrm{HS}}(\ell_{2}X)$ and $\Xi:=(\Xi_{n})_{n}\in\mathcal{H}\subset\prod_{\mathcal{U}}\operatorname{\mathrm{HS}}(\ell_{2}X)$ are unit vectors. By the assumption of exactness, $\pi$ drops to the representation of $A:=\pi(\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}\left\lvert\Gamma\curvearrowright(X\times\mathbb{N})\right\rvert)\otimes{\operatorname{\mathrm{C}^{*}}(\lambda_{X}(\Gamma))}^{\mathrm{op}}$, where we simply write $\pi$ as its restriction to $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}\left\lvert\Gamma\curvearrowright(X\times\mathbb{N})\right\rvert\otimes 1$.
We invoke Hahn-Banach theorem.
Lemma 8.
Let $A\subset\mathbb{B}(\mathcal{H})$ be a $\operatorname{\mathrm{C}^{*}}$-subalgebra. Then the set of vector states $\varphi_{\zeta}$ for $\zeta\in\mathcal{H}\otimes\ell_{2}$ is dense in the state space of $A$ w.r.t. weak topology.
By Hahn-Banach theorem for $A\subset\mathbb{B}(\mathcal{H}\otimes\overline{\ell_{2}X})$, there is a net of unit vectors $\zeta^{i}\in\mathcal{H}\otimes\overline{\ell_{2}X}\otimes\ell_{2}$ such that $\varphi_{\zeta^{i}}\to\varphi_{\Xi}$ in weak topology of $A^{*}$. Denote $\mathcal{K}:=\overline{\ell_{2}X}\otimes\ell_{2}$. We may assume that $\zeta^{i}$ are in the algebraic tensor product $\mathcal{H}_{0}\odot\mathcal{K}$, take a lift $(\zeta^{i}_{n})_{n}$ of $\zeta^{i}$ in $\prod_{n}\operatorname{\mathrm{HS}}(\ell_{2}X)\odot\mathcal{K}$, and define unit vectors $\eta^{i}\in\mathcal{H}_{0}$ by $\eta^{i}_{n}(x,y):=\left\lVert\zeta^{i}_{n}(x,y)\right\rVert_{\mathcal{K}}$, where $\zeta^{i}_{n}(x,y)\in\mathcal{K}$ is the $(x,y)$-entry of $\zeta^{i}_{n}$.
Let us now sort out what vectors define the vector states on which algebras. $\Xi\in\mathcal{H}$ and $\zeta^{i}\in\mathcal{H}\otimes\mathcal{K}$ define states on $A$, $\xi_{n}\in\ell_{2}X$ and $\eta^{i}\in\mathcal{H}$ define states on $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}\left\lvert\Gamma\curvearrowright(X\times\mathbb{N})\right\rvert$ by taking the ultralimit.
Lemma 9.
For $\varepsilon>0,\ \gamma\in\Gamma,\ f\in\ell_{\infty}(X\times\mathbb{N})\subset\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}\left\lvert\Gamma\curvearrowright(X\times\mathbb{N})\right\rvert$ and sufficient large $i$, the following hold.
  • $\mathopen{\langle}\eta^{i},h^{\gamma}\gamma\eta^{i}\mathclose{\rangle}=\varphi_{\eta^{i}}(h^{\gamma}\gamma)\geq\left\lvert\varphi_{\zeta^{i}}(h^{\gamma}\gamma\otimes{\gamma^{-1}}^{\mathrm{op}})\right\rvert\approx_{\varepsilon}\varphi_{\Xi}(h^{\gamma}\gamma\otimes{\gamma^{-1}}^{\mathrm{op}})=1$.
  • $\left\lVert h^{\gamma}\gamma\eta^{i}\right\rVert^{2}=\varphi_{\eta^{i}}(\left\lvert h^{\gamma}\gamma\right\rvert^{2})=\varphi_{\zeta^{i}}(\left\lvert h^{\gamma}\gamma\otimes{\gamma^{-1}}^{\mathrm{op}}\right\rvert^{2})\approx_{\varepsilon}\varphi_{\Xi}(\left\lvert h^{\gamma}\gamma\otimes{\gamma^{-1}}^{\mathrm{op}}\right\rvert^{2})=1$.
  • $\varphi_{\eta^{i}}(\left\lvert f\right\rvert^{2})=\varphi_{\zeta^{i}}(\left\lvert f\right\rvert^{2})\approx_{\varepsilon}\varphi_{\Xi}(\left\lvert f\right\rvert^{2})=\lim_{\mathcal{U}}\varphi_{\xi_{n}}(\left\lvert f_{n}\right\rvert^{2})$.
Proof.
Note that $(h_{\gamma}\gamma\otimes{\gamma^{-1}}^{\mathrm{op}})\Xi=h_{\gamma}\gamma\Xi\gamma^{-1}=\Xi$ and $\left\lVert f_{n}\Xi_{n}\right\rVert^{2}=\left\lVert f_{n}\xi_{n}\right\rVert^{2}$. One has
$\mathopen{\langle}\eta^{i},h^{\gamma}\gamma\eta^{i}\mathclose{\rangle}$ $=\lim_{\mathcal{U}}\sum_{x,y}h^{\gamma}_{n}(x)\eta^{i}_{n}(x,y)\eta^{i}_{n}(\gamma^{-1}x,y)$
$\geq\lim_{\mathcal{U}}\left\lvert\sum_{x,y}h^{\gamma}_{n}(x)\mathopen{\langle}\zeta^{i}_{n}(x,y),\overline{\gamma}\zeta^{i}_{n}(\gamma^{-1}x,y)\mathclose{\rangle}_{\mathcal{K}}\right\rvert$
$=\left\lvert\mathopen{\langle}\zeta^{i},(h^{\gamma}\gamma\otimes{\gamma^{-1}}^{\mathrm{op}})\zeta^{i}\mathclose{\rangle}\right\rvert,$
$\left\lVert h^{\gamma}\gamma\eta^{i}\right\rVert^{2}$ $=\lim_{\mathcal{U}}\sum_{x,y}h^{\gamma}_{n}(x)^{2}\eta^{i}_{n}(\gamma^{-1}x,y)^{2}$
$=\lim_{\mathcal{U}}\sum_{x,y}h^{\gamma}_{n}(x)^{2}\left\lVert\overline{\gamma}\zeta^{i}_{n}(\gamma^{-1}x,y)\right\rVert^{2}_{\mathcal{K}}$
$=\left\lVert(h_{\gamma}\gamma\otimes{\gamma^{-1}}^{\mathrm{op}})\zeta^{i}\right\rVert^{2},$
$\left\lVert f\eta^{i}\right\rVert^{2}$ $=\lim_{\mathcal{U}}\sum_{x,y}\left\lvert f_{n}(x)\right\rvert^{2}\eta^{i}_{n}(x,y)^{2}$
$=\lim_{\mathcal{U}}\sum_{x,y}\left\lvert f_{n}(x)\right\rvert^{2}\left\lVert\zeta^{i}_{n}(x,y)\right\rVert^{2}_{\mathcal{K}}$
$=\left\lVert f\zeta^{i}\right\rVert^{2},$
and $\varphi_{\zeta^{i}}(x)\approx_{\varepsilon}\varphi_{\Xi}(x)$ for $x\in A$ and suffcient large $i$. ∎
Thus, $h^{\gamma}\gamma\eta^{i}\approx_{2\sqrt{\epsilon}}\eta^{i}$ and $\varphi_{\eta^{i}}(f)\approx_{\varepsilon}\lim_{\mathcal{U}}\varphi_{\xi_{n}}(f_{n})$ for $f\in\ell_{\infty}(X\times\mathbb{N})$. One has
$\varphi_{\eta^{i}}(fh^{\gamma}\gamma)$ $=\mathopen{\langle}\eta^{i},fh^{\gamma}\gamma\eta^{i}\mathclose{\rangle}$
$\approx_{2\sqrt{\epsilon}}\mathopen{\langle}\eta^{i},f\eta^{i}\mathclose{\rangle}$
$\approx_{\varepsilon}\lim_{\mathcal{U}}\mathopen{\langle}\xi_{n},f_{n}\xi_{n}\mathclose{\rangle}$
$=\lim_{\mathcal{U}}\mathopen{\langle}\xi_{n},f_{n}h^{\gamma}_{n}\gamma\xi_{n}\mathclose{\rangle}$
$=\lim_{\mathcal{U}}\phi_{\xi_{n}}(f_{n}h^{\gamma}_{n}\gamma).$
Therefore, $\varphi_{\eta^{i}}\to\lim_{\mathcal{U}}\varphi_{\xi_{n}}$ in weak topology of $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}\left\lvert\Gamma\curvearrowright(X\times\mathbb{N})\right\rvert^{*}$.
Lemma 10.
For every unit vector $\eta\in\mathcal{H}_{0}$, there exists $S>0$ such that $\left\lvert\varphi_{\eta}(a)\right\rvert\leq\lim_{\mathcal{U}}\left\lVert\Phi_{S}(a_{n})\right\rVert$ for every $a\in\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}\left\lvert\Gamma\curvearrowright(X\times\mathbb{N})\right\rvert$.
This lemma is essentially observed in [Sak20, Theorem 3.17].
Proof.
Define $\omega^{y}_{n}\in\ell_{2}\operatorname{Ball}(y,S)$ as $\omega^{y}_{n}(x):=\eta_{n}(x,y)$ and $\omega_{n}:=\oplus_{y}\omega^{y}_{n}\in\bigoplus_{y}\ell_{2}\operatorname{Ball}(y,S)$.
$\varphi_{\eta}(x)=\lim_{\mathcal{U}}\mathopen{\langle}\eta_{n},x_{n}\eta_{n}\mathclose{\rangle}=\lim_{\mathcal{U}}\sum_{y}\mathopen{\langle}\omega^{y}_{n},x_{n}\omega^{y}_{n}\mathclose{\rangle}=\lim_{\mathcal{U}}\mathopen{\langle}\omega_{n},\Phi_{S}(x_{n})\omega_{n}\mathclose{\rangle}$
shows that $\varphi_{\eta}$ factors through $\Phi_{S}$. Note that $\sum_{y}\left\lVert\omega^{y}_{n}\right\rVert^{2}=1$. ∎
Recall $b_{n}$ was taken as $\lim_{\mathcal{U}}\varphi_{\xi_{n}}(b_{n})\approx_{\varepsilon}1$, $\lim_{\mathcal{U}}\left\lVert\Phi_{S}(b_{n})\right\rVert\leq 0.1$ for every $S>0$ and $b_{n}\in\mathbb{C}^{\leq R}_{\mathrm{u}}(X)$ for some $R>0$. Therefore, $b:=\oplus_{n}b_{n}\in\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}\left\lvert\Gamma\curvearrowright(X\times\mathbb{N})\right\rvert$ gives a contradiction. We obtain the main theorem.
We note that we only need min-continuity of $\pi:\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}\left\lvert\Gamma\curvearrowright X\right\rvert\otimes{\operatorname{\mathrm{C}^{*}}(\lambda_{X}(\Gamma))}^{\mathrm{op}}\to\mathbb{B}(\mathcal{H})$. Since $\xi_{n}$ has exponential decay and sparse approximate supports, we define $h^{\gamma}\in\ell_{\infty}X$ as $h^{\gamma}=h^{\gamma}_{n}$ on a sufficiently large neighborhood of the approximate support of $\xi_{n}$.
4. Appendix
The purpose of this appendix is to show that every ulf coarse space comes from some group action up to coarse equivalence.
Definition 11 (coarse space).
$\mathcal{E}\subset 2^{X\times X}$ is called a coarse structure on a set $X$ when $\mathcal{E}$ forms ideal by the inclusion order (i.e., lower and directed set) and satisfies the following group-like axioms:
  • $\Delta_{X}:=\{(x,x)\in X\times X\ \mid\ x\in X\}\in\mathcal{E}$;
  • $E\circ F:=\{(x,z)\in X\times X\ \mid\ (x,y)\in E,(y,z)\in F\}\in\mathcal{E}$ for every $E,F\in\mathcal{E}$;
  • $E^{-1}:=\{(y,x)\in X\times X\ \mid\ (x,y)\in E\}\in\mathcal{E}$ for every $E\in\mathcal{E}$.
Example 12 (extended metric space, graph).
Let $(X,d)$ be a metric space (possiblly $d$ takes a value in $[0,\infty]$). We associate the metric coarse structure
:=$\mathcal{E}:=\{E\subset X\times X\ \mid\ d\ \text{is bounded on}\ E\}.$
Especially, when $d$ comes from the graph metric of a graph $(X,E)$, this coincides with the coarse structure generated by $E$, i.e. the smallest coarse structure containing $E$.
We note that a coarse space $(X,\mathcal{E})$ comes from a graph if and only if $\mathcal{E}$ is finitely generated (actually, singly generated), and $(X,\mathcal{E})$ comes from an extended metric if and only if $\mathcal{E}$ is countably generated. To see this, take a generator set $\{E_{n}\}_{n=0}^{\infty}$ as $E_{0}=\Delta_{X},\ E_{n}\circ E_{m}\subset E_{n+m},\ E^{-1}_{n}=E_{n}$, and define $d(x,y):=\min\{\,n\mid(x,y)\in E_{n}\,\}$. We define the Schreier graph (coarse space) for a general discrete group and its action.
Definition 13 (Schreier graph).
For a discrete group action $\Gamma\curvearrowright X$, we define the Schreier graph $\left\lvert\Gamma\curvearrowright X\right\rvert$ as the ulf coarse structure generated by $\{\operatorname{graph}\gamma\}_{\gamma\in\Gamma}$, where $\operatorname{graph}\gamma:=\{\,(\gamma x,x)\in X\times X\mid x\in X\,\}$
Definition 14 (ulf).
For a coarse space $(X,\mathcal{E})$, we assume that $\sup_{x}\left\lvert S\cap(X\times\{x\})\right\rvert<\infty$ for every $S\in\mathcal{E}$, which is called uniformly locally finite, or ulf in short.
Definition 15 (uniform Roe algebra).
$\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X,\mathcal{E}):=\overline{\bigcup_{S\in\mathcal{E}}\{\,a\in\mathbb{B}(\ell_{2}X)\mid a(x,y)=0\ \text{if}\ (x,y)\notin S\,\}}$
extends our definition of the uniform Roe algebra for graphs to general coarse spaces.
Lemma 16 (coloring lemma).
Let $S\subset X\times X$ be that $d:=\sup_{x\in X}\left\lvert S\cap(\{x\}\times X)\right\rvert+\left\lvert S\cap(X\times\{x\})\right\rvert<\infty$. There are partial translations $\{T_{i}\}_{i=1}^{d}$ such that $S$ is disjoint union of $\operatorname{graph}T_{i}$.
Recall that a partial translation on $X$ is an injection from some subset of $X$ to $X$. We rewrite the proof in Lemma 2. of [Oza23] in our setting.
Proof.
Consider partial translations $\{T_{i}\}_{i=1}^{d}$ such that
$\bigsqcup_{i=1}^{d}\operatorname{graph}T_{i}\subset S$
and take a maximal (w.r.t. the inclusion order in $X\times X$) one. We claim that the above inclusion is indeed equal. Suppose not and take $(x,y)$ from the set difference. By the maximality, every $\operatorname{graph}T_{i}$ should meet with $(S\setminus\{(x,y)\})\cap(\{x\}\times X\cup X\times\{y\})$, but this set has the cardinality less than $d$ because $(x,y)\in S$. The pigeonhole principle gives a contradiction. ∎
Remark 17.
Every ulf coarse space is realized as the Schreier graph of (possibly infinite generated) free group action up to coarse equivalence. When the coarse space comes from a graph (i.e., finitely generated as a coarse space), the free group can be taken as a finitely generated one.
Proof.
Since ulf coarse structure is generated by partial translations, the $2\times 2$ trick (namely, a partial isometry dilates to a self-adjoint unitary) works well:
$\mathrm{U}(v):=\begin{pmatrix}1-vv^{*}&v\\ v^{*}&1-v^{*}v\end{pmatrix}$
has $\{0,1\}$-valued matrix entries when $v$ is a partial translation, and gives an involution on $X\sqcup X$. These involutions for every partial translation (including $v=1$) generate the coarse structure on $X\sqcup X$, the graph Cartesian product of $X$ and $K_{2}$.
Suppose that $X$ has a finite generator set $S$ of partial translations. Then, we use the above $2\times 2$ trick for $S\cup\{1\}$, and one sees that the resultant coarse structure on $X\sqcup X$ is the Cartesian product of $X$ and $K_{2}$. This is coarsely equivalent to $X$. ∎
By this proof, our main theorem also holds in the partial translation setting.
Remark 18.
Let $\Gamma$ be a family of partial translations on a set $X$. Then, $\operatorname{\mathrm{C}^{*}}(\Gamma)\subset\mathbb{B}(\ell_{2}X)$ is exact, if and only if $\Gamma$ generates the coarse structure $\mathcal{E}=\mathopen{\langle}\Gamma\mathclose{\rangle}$ on $X$ with property A.
Proof.
Since $\operatorname{\mathrm{C}^{*}}(\Gamma)\subset\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X,\mathcal{E})$, it suffices to show that exactness of $\operatorname{\mathrm{C}^{*}}(\Gamma)$ implies nuclearity of $\operatorname{\mathrm{C}^{*}_{\mathrm{u}}}(X,\mathcal{E})$. Let $\Lambda$ be the group generated by involutions on $X\sqcup X$, obtained by the above construction for $\Gamma\cup\{1\}$. Then $\operatorname{\mathrm{C}^{*}}(\lambda_{X\sqcup X}(\Lambda))\subset\mathbb{M}_{2}\otimes\widetilde{\operatorname{\mathrm{C}^{*}}(\Gamma)}$, where $\widetilde{A}$ stands the unitization of $A$. By passing exactness to $\operatorname{\mathrm{C}^{*}}(\lambda_{X\sqcup X}(\Lambda))$, $\left\lvert\Lambda\curvearrowright(X\sqcup X)\right\rvert$ and $(X,\mathcal{E})$ is property A. ∎